# No evidence for $H_{0}$ anisotropy from Tully–Fisher or supernova distances

### v1

@@ Abstract
Claims of anisotropy in the Hubble constant have been made based on direct distance tracers such as Tully–Fisher galaxies and Type Ia supernovae. We revisit these using the CosmicFlows-4 Tully–Fisher W1 subsample, 2MTF and SFI++ Tully–Fisher catalogues, and the Pantheon+ supernova compilation (restricted to $z<0.05$), including a dipole in either the Tully–Fisher zero-point or the standardised supernova absolute magnitude. Our forward-modelling framework jointly calibrates the distance relation, marginalises over distances, and accounts for peculiar velocities using a linear-theory reconstruction. We compare the anisotropic and isotropic model using the Bayesian evidence. In the CosmicFlows-4 sample, we infer a zero-point dipole of amplitude $0.087\pm 0.019$ mag, or $4.1\pm 0.9$ per cent when expressed as a dipole in the Hubble parameter. This is consistent with previous estimates but at higher significance: model comparison yields odds of 877:1877\!:\!1 in favour of including the zero-point dipole. In Pantheon+ we infer zero-point dipole amplitude of $0.049\pm 0.013$ mag, or $2.3\pm 0.6$ per cent when expressed as a dipole in the Hubble parameter. However, by allowing for a radially varying velocity dipole, we show that the anisotropic model captures local flow features (or possibly systematics) in the data rather than an actual linearly growing effective bulk flow caused by anisotropy in the zero-point or expansion rate. Inferring a more general bulk flow curve we find results fully consistent with expectations from the standard cosmological model.
@@
pubyear: 2025pagerange: No evidence for $H_{0}$ anisotropy from Tully–Fisher or supernova distances9
@@ 1 Introduction
The cosmological principle—that the Universe is isotropic and homogeneous on sufficiently large scales—plays a foundational role in modern cosmology. Combined with the assumption of General Relativity as the theory of gravity, it leads to the Friedmann–Robertson–Walker metric which underpins the concordance model of cosmology, $\Lambda$-cold dark matter. It is therefore crucial to test if the cosmological principle is satisfied and hence fit for the purpose of providing such a bedrock. Recent years have seen increased scrutiny of the cosmological principle, for example in the cosmic microwave background (cosmic microwave background; Jones et al. 2023; Gaztañaga & Sravan Kumar 2024), Type Ia supernovae (Kalbouneh et al., 2023; Hu et al., 2024; Horstmann et al., 2022; Sorrenti et al., 2023a; Sah et al., 2024; Mc Conville & Ó Colgáin, 2023; Krishnan et al., 2022; Zhai & Percival, 2022; Rahman et al., 2022; Cowell et al., 2023; Sorrenti et al., 2025), direct distance tracers (Watkins et al., 2023; Boubel et al., 2025), galaxy clusters (Migkas et al., 2021; Pandya et al., 2024) and distant radio galaxies, quasars and gamma ray bursts (Secrest et al., 2022a; Dam et al., 2023; Secrest et al., 2025; Luongo et al., 2022) (for a review, see Aluri et al. 2023). Any evidence for anisotropy must be carefully validated before violation of so fundamental a principle as the cosmological principle can be concluded.
@@
One powerful test of the cosmological principle involves examining the local expansion rate $H_{0}$ in different directions, which can be achieved by using direct distance tracers such as Tully–Fisher relation galaxies and Type Ia supernovae. This is doubly important as it may also be relevant for the Hubble tension, which is derived assuming isotropic expansion (Riess et al., 2022; Di Valentino et al., 2025). Without calibration through a lower rung of the distance ladder, an anisotropy in $H_{0}$ is completely degenerate with an anisotropy in the normalisation of the TFR and the absolute magnitude of supernovae. While it is therefore not possible to assess an $H_{0}$ anisotropy directly with this method, it seems implausible that a null detection of the degenerate combination would arise from compensated anisotropies in both $H_{0}$ and the normalisation of the standardising relation.
@@
In this paper, we revisit claims of anisotropy in $H_{0}$ by fully forward-modelling the Tully–Fisher relation and supernova observables. Our starting point is the recent study of Boubel et al. (2025) (hereafter B25), who conclude weak evidence for anisotropy in the CosmicFlows-4 W1 Tully–Fisher relation subsample using a partial forward model. Following B25, we adopt the first-order deviation from isotropy in spherical harmonics, namely a dipolar modulation to the degenerate combination of the Tully–Fisher relation or supernova normalisation and $H_{0}$. Our approach jointly calibrates the distance relation, marginalizes over distances and latent parameters describing the true values of the observables, and corrects for peculiar velocities using a linear-theory reconstruction based on Carrick et al. (2015) (hereafter C15). We generalise the analysis of B25 to three further independent data sets to provide a more comprehensive and robust assessment of local anisotropy, namely the 2MTF and SFI++ Tully–Fisher relation catalogues and the Pantheon+ supernova compilation (restricted to $z<0.05$). We also, for the first time, explicitly investigate the impact of dust extinction on the results by investigating three qualitatively different maps with different priors on the extinction coefficients. This is potentially an important systematic because extinction is anisotropic across the sky.
@@
However, we caution that although we search for a cosmological principle-violating anomaly, this is done under the assumption of the C15 reconstruction to account for peculiar velocities, which itself assumes the cosmological principle. In principle, the C15 reconstruction could already contain part of a cosmological dipole, which would then be subtracted in our analysis. Nevertheless, the bulk flow in C15 shows no significant deviation from $\Lambda$-cold dark matter (see also e.g. Boruah et al. 2020; Stiskalek et al. 2025a). Thus, the results presented here should be interpreted as a dipole in $H_{0}$ that would be superimposed on the dipole already present in the C15 peculiar velocity field, which is not cosmological principle-violating.
@@
The remainder of this paper is structured as follows. In Section˜2, we describe the data sets used in our analysis, including the CosmicFlows-4 Tully–Fisher relation sample, the 2MTF sample, the SFI++ catalogue, and the Pantheon+ supernova compilation. In Section˜3, we present our methodology, including the forward-modelling framework, the peculiar velocity modelling scheme, the Bayesian evidence calculation and the mock data procedure. Section˜4 presents the results on both mock and observed data, while Section˜5 presents further comparison with B25, discusses the more general ramifications of our results, and concludes. Appendix A describes the flow model in full, and Appendix B shows the complete posterior distribution for the CosmicFlows-4 W1 Tully–Fisher relation inference.
@@
All logarithms are base-10 unless otherwise stated. We use the notation $\mathcal{N}(x;\mu,\sigma)$ to denote a one-dimensional normal distribution with mean $\mu$ and standard deviation $\sigma$, evaluated at $x$; in higher dimensions $\mu$ is a vector and $\sigma$ is replaced by a covariance matrix. We define $h\equiv H_{0}/\left(100\mathrm{km}\,\mathrm{s}^{-1}\,\mathrm{Mpc}^{-1}\right)$, where $H_{0}$ is the Hubble constant.
@@ 2 Data
To test the hypothesis of a dipole in the zero-point calibration of redshift-independent distance indicators, or, equivalently, in the Hubble constant, we analyse multiple low-redshift catalogues. Our primary data set is the Tully–Fisher relation subsample of the CosmicFlows-4 catalogue (CosmicFlows-4; Tully et al. 2023), restricted to photometry in the all-sky W1 band. In addition, we consider two other all-sky Tully–Fisher relation samples: 2MTF (Masters et al., 2008; Hong et al., 2019) and the SFI++ catalogue (Springob et al., 2007). We also analyse the Pantheon+ compilation of Type Ia supernova (Scolnic et al., 2022), along with its reanalysis by Lane et al. (2024).
Since any large-scale dipole in the zero-point is degenerate with the local peculiar velocity field, we account for peculiar velocities using the linear reconstruction of C15. This choice, widely adopted in the literature (e.g. Boruah et al., 2020; Said et al., 2020; Carr et al., 2022; Boubel et al., 2024b, 2025), allows for direct comparison with recent studies, particularly B25, and facilitates consistent treatment of large-scale flows across all samples.
@@ 2.1 Tully–Fisher samples
Our first method for obtaining redshift-independent distance estimates is the Tully–Fisher relation (Tully–Fisher relation; Tully & Fisher 1977), which relates a spiral galaxy’s rotational velocity to its absolute magnitude $M$. Given an observed apparent magnitude, this relation yields a distance modulus and thereby a peculiar velocity, since the observed redshift is a function of the distance and peculiar velocity. The linewidth parameter $\eta$ is defined as
@@ 2.1.1 CosmicFlows-4 TFR sample
We use the CosmicFlows-4 Tully–Fisher relation catalogue, a subset of the broader CosmicFlows-4 compilation (Tully et al., 2023), consisting of $9792$ galaxies with $z_{\rm CMB}\lesssim 0.05$ and no strict apparent magnitude limit (Kourkchi et al., 2020a, b). Our analysis uses photometry exclusively in the all-sky Wide-field Infrared Survey Explorer W1 band (same as B25), with additional selection criteria applied. We require $\eta>-0.3$ (to eliminate dwarf and low-mass galaxies, which may follow a different Tully–Fisher relation or have a higher scatter), Galactic latitude $|b|>7.5^{\circ}$ to exclude the Galactic Zone of Avoidance, and a photometric quality flag of $5$ (“best”). After these cuts, the final W1 sample contains $3246$ galaxies. The highest photometric quality requirement excludes $\sim$$1300$ galaxies from the W1 subsample, while the Galactic latitude cut removes only 44 galaxies from the final sample. The Zone of Avoidance is typically defined as $|b|<5^{\circ}$ (e.g. Staveley-Smith et al. 1998), making our choice relatively conservative. In contrast, B25 do not apply either cut. Since the publicly released CosmicFlows-4 catalogue does not provide magnitude uncertainties, we adopt a conservative fiducial uncertainty of $0.05~\mathrm{mag}$, following Kourkchi et al. (2019). However, this uncertainty is subdominant to the intrinsic scatter in the Tully–Fisher relation, which is approximately $0.35~\mathrm{mag}$ (see Fig.˜2 or Fig.˜9). Moreover, unlike B25, who impose a lower redshift cut of $cz_{\rm CMB}>3000~\mathrm{km}\,\mathrm{s}^{-1}$, we do not apply a lower redshift limit in our main analysis, but discuss its impact in Section˜5.
@@
Next, we consider the 2MASS Tully–Fisher Survey (2MTF), an all-sky sample of 2,062 spiral galaxies with Tully–Fisher relation measurements extending to redshifts of $z_{\mathrm{CMB}}\lesssim 0.03$ (Masters et al., 2008; Hong et al., 2019). The survey is selected in the $K$ band with an apparent magnitude limit of $K<11.25$. We use the version of the catalogue compiled by Boruah et al. (2020), which removes duplicates from the SFI++ sample, includes only $K$-band photometry, and applies a quality cut on linewidths, retaining only galaxies with $0.1<\eta<0.2$. These cuts leave $1247$ galaxies.
@@ 2.1.3 SFI++ TFR sample
We also consider the SFI++ catalogue, an all-sky sample comprising galaxies and groups with TFR measurements extending to redshifts of $z_{\mathrm{CMB}}\lesssim 0.05$ (Masters et al., 2006; Springob et al., 2007). Unlike 2MTF, the survey does not impose a strict apparent magnitude limit, and photometry is provided in the $I$ band. In this work, we use the galaxy-only version of the catalogue compiled by Boruah et al. (2020), who apply quality cuts to select galaxies within the 2M++ footprint, impose a lower linewidth threshold to exclude low-mass galaxies, and use an iterative procedure to reject Tully–Fisher relation outliers. While Boruah et al. (2020) also impose a strict upper linewidth selection to ensure linearity of the TFR, we relax this constraint and allow for Tully–Fisher relation curvature, resulting in a final sample of $2010$ galaxies.
@@ 2.2 Pantheon+ supernovae sample
Type Ia supernovae are widely used as standardisable candles in cosmology. The SALT2 model standardises their light curve (Guy et al., 2007), yielding a standardised apparent magnitude via the Tripp formula (Tripp, 1998):
@@
The Pantheon+ data set is a compilation of $1701$ spectroscopically confirmed Type Ia supernovae spanning redshifts from $z\sim 0.001$ to $\sim 2.3$ (Scolnic et al., 2022; Brout et al., 2022; Peterson et al., 2022; Carr et al., 2022). For consistency with our peculiar velocity modelling and to match the redshift range of the CosmicFlows-4 sample, we restrict the data to $z_{\rm CMB}\leq 0.05$, yielding a subset of $525$ supernovae. Pantheon+ combines the original Pantheon sample (Scolnic et al., 2018) with updated low- and high-redshift supernovae, incorporating improved photometric calibration, light-curve standardisation, and systematic uncertainty modelling. Distance moduli are derived using the SALT2 light-curve fitter and corrected for selection effects via the BEAMS with Bias Corrections (BBC) method (Kessler & Scolnic, 2017), which introduces an additive bias correction term in Eq.˜3. We use such bias-corrected apparent magnitudes $m_{\rm corr}$ from the Pantheon+ sample, which include a fiducial Tripp calibration. We therefore sample only $M_{\rm SN}$, not the light curve stretch and colour coefficients of the Tripp calibration.
@@
A reanalysis of the Pantheon+ data was presented by Lane et al. (2024), who constructed a supernova covariance matrix designed to minimise dependence on the assumed cosmology and peculiar velocities. Unlike the standard Pantheon+ covariance matrix and bias-corrected apparent magnitudes, which assume a fiducial cosmology and Tripp parameters, the Lane et al. covariance matrix does not include any contribution from the Tripp parametrisation. This enables simultaneous inference of the Tripp coefficients, with the standardised apparent magnitudes expressed as in Eq.˜3, and no further bias corrections. We treat this as a variant of the Pantheon+ data, denoted “Pantheon+Lane” henceforth.

### v2

# No evidence for local $H_{0}$ anisotropy from Tully–Fisher or supernova distances

@@ Abstract
Claims of local ($z\lesssim 0.05$) anisotropy in the Hubble constant have been made based on direct distance tracers such as Tully–Fisher galaxies and Type Ia supernovae. We revisit these using the CosmicFlows-4 Tully–Fisher W1 subsample, 2MTF and SFI++ Tully–Fisher catalogues, and the Pantheon+ supernova compilation (all restricted to $z<0.05$), including a dipole in either the Tully–Fisher zero-point or the standardised supernova absolute magnitude. Our forward-modelling framework jointly calibrates the distance relation, marginalises over distances, and accounts for peculiar velocities using a linear-theory reconstruction. We compare the anisotropic and isotropic model using the Bayesian evidence. In the CosmicFlows-4 sample, we infer a zero-point dipole of amplitude $0.087\pm 0.019$ mag, or $4.1\pm 0.9$ per cent when expressed as a dipole in the Hubble parameter. This is consistent with previous estimates but at higher significance: model comparison yields odds of 877:1877\!:\!1 in favour of including the zero-point dipole. In Pantheon+ we infer zero-point dipole amplitude of $0.049\pm 0.013$ mag, or $2.3\pm 0.6$ per cent when expressed as a dipole in the Hubble parameter. However, by allowing for a radially varying velocity dipole, we show that the anisotropic zero-point model captures local flow features (or possibly systematics) in the data rather than an actual linearly growing effective bulk flow caused by anisotropy in the zero-point or expansion rate. Crucially, inferring a more general bulk flow curve we find results fully consistent with expectations from the standard cosmological model.
@@
pubyear: 2025pagerange: No evidence for local $H_{0}$ anisotropy from Tully–Fisher or supernova distances9
@@ 1 Introduction
The cosmological principle—that the Universe is isotropic and homogeneous on sufficiently large scales—plays a foundational role in modern cosmology. Combined with the assumption of General Relativity as the theory of gravity, it leads to the Friedmann–Robertson–Walker metric which underpins the concordance model of cosmology, $\Lambda$-cold dark matter. It is therefore crucial to test if the cosmological principle is satisfied and hence fit for the purpose of providing such a bedrock. Recent years have seen increased scrutiny of the cosmological principle, for example in the cosmic microwave background (cosmic microwave background; CMB_anisotropy; Sravan), Type Ia supernovae (Basheer; Hu; SN_1; SN_2; Sah; Conville; Krishnan; Zhai; Rahman_2022; Cowell_2023; Sorrenti_2025), direct distance tracers (Watkins_2023; Boubel_2025), galaxy clusters (Migkas_2021; Pandya_2024) and distant radio galaxies, quasars and gamma ray bursts (EB_1; EB_2; colloquium; Luongo_2022) (for a review, see CP_review). Moreover, increasing attention has been directed toward defining “expansion” in metrics beyond Friedmann–Robertson–Walker to enable model-independent tests (Maartens_2024; Kalbouneh_2024; Kalbouneh_2025; Sarma_2025). Any evidence for anisotropy must be carefully validated before violation of so fundamental a principle as the cosmological principle can be concluded.
@@
One powerful test of the cosmological principle involves examining the local expansion rate $H_{0}$ in different directions, which can be achieved by using direct distance tracers such as Tully–Fisher relation galaxies and Type Ia supernovae. This is doubly important as it may also be relevant for the Hubble tension, which is derived assuming isotropic expansion (Riess_2022; cosmoverse). Without calibration through a lower rung of the distance ladder, an anisotropy in $H_{0}$ is completely degenerate with an anisotropy in the normalisation of the TFR and the absolute magnitude of supernovae. While it is therefore not possible to assess an $H_{0}$ anisotropy directly with this method, it seems implausible that a null detection of the degenerate combination would arise from compensated anisotropies in both $H_{0}$ and the normalisation of the standardising relation.
@@
In this paper, we revisit claims of anisotropy in $H_{0}$ by fully forward-modelling the Tully–Fisher relation and supernova observables at $z<0.05$. Our starting point is the recent study of Boubel_2025 (hereafter Boubel_2025), who conclude weak evidence for anisotropy in the CosmicFlows-4 W1 Tully–Fisher relation subsample using a partial forward model. Following Boubel_2025, we adopt the first-order deviation from isotropy in spherical harmonics, namely a dipolar modulation to the degenerate combination of the Tully–Fisher relation or supernova normalisation and $H_{0}$. Our approach jointly calibrates the distance relation, marginalizes over distances and latent parameters describing the true values of the observables, and corrects for peculiar velocities using a linear-theory reconstruction based on Carrick_2015 (hereafter Carrick_2015). We generalise the analysis of Boubel_2025 to three further independent data sets to provide a more comprehensive and robust assessment of local anisotropy, namely the 2MTF and SFI++ Tully–Fisher relation catalogues and the Pantheon+ supernova compilation (restricted to $z<0.05$). We also, for the first time, explicitly investigate the impact of dust extinction on the results by investigating three qualitatively different maps with different priors on the extinction coefficients. This is potentially an important systematic because extinction is anisotropic across the sky.
@@
However, we caution that although we search for a cosmological principle-violating anomaly, this is done under the assumption of the Carrick_2015 reconstruction to account for peculiar velocities, which itself assumes the cosmological principle. In principle, the Carrick_2015 reconstruction could already contain part of a cosmological dipole, which would then be subtracted in our analysis. Nevertheless, the bulk flow in Carrick_2015 shows no significant deviation from $\Lambda$-cold dark matter (see also e.g. Boruah_2019; VF_olympics). Thus, the results presented here should be interpreted as a dipole in $H_{0}$ that would be superimposed on the dipole already present in the Carrick_2015 peculiar velocity field, which is not cosmological principle-violating. Nevertheless, we also consider the total bulk flow curve from both Carrick_2015 and a superimposed radially varying velocity dipole to test whether the resulting flow is in tension with $\Lambda$-cold dark matter expectations.
@@
The remainder of this paper is structured as follows. In Section˜2, we describe the data sets used in our analysis, including the CosmicFlows-4 Tully–Fisher relation sample, the 2MTF sample, the SFI++ catalogue, and the Pantheon+ supernova compilation. In Section˜3, we present our methodology, including the forward-modelling framework, the peculiar velocity modelling scheme, the Bayesian evidence calculation and the mock data procedure. Section˜4 presents the results on both mock and observed data, while Section˜5 presents further comparison with Boubel_2025, discusses the more general ramifications of our results, and concludes. Appendix A describes the flow model in full, and Appendix B shows the complete posterior distribution for the CosmicFlows-4 W1 Tully–Fisher relation inference.
@@
All logarithms are base-10 unless otherwise stated. We use the notation $\mathcal{N}(x;\mu,\sigma)$ to denote a one-dimensional normal distribution with mean $\mu$ and standard deviation $\sigma$, evaluated at $x$; in higher dimensions $\mu$ is a vector and $\sigma$ is replaced by a covariance matrix. We define $h\equiv H_{0}/\left(100~\mathrm{km}\,\mathrm{s}^{-1}\,\mathrm{Mpc}^{-1}\right)$, where $H_{0}$ is the Hubble constant.
@@ 2 Data
To test the hypothesis of a dipole in the zero-point calibration of redshift-independent distance indicators, or, equivalently, in the Hubble constant, we analyse multiple low-redshift catalogues. Our primary data set is the Tully–Fisher relation subsample of the CosmicFlows-4 catalogue (CosmicFlows-4; Tully_2023), restricted to photometry in the all-sky W1 band. In addition, we consider two other all-sky Tully–Fisher relation samples: 2MTF (Masters_2008; Hong_2019) and the SFI++ catalogue (Springob_2007). We also analyse the Pantheon+ compilation of Type Ia supernova (Scolnic_2022), along with its reanalysis by Lane_2024.
Since any large-scale dipole in the zero-point is degenerate with the local peculiar velocity field, we account for peculiar velocities using the linear reconstruction of Carrick_2015. This choice, widely adopted in the literature (e.g. Boruah_2019; Said_2020; Carr_2022; Boubel_2024_H0; Boubel_2025), allows for direct comparison with recent studies, particularly Boubel_2025, and facilitates consistent treatment of large-scale flows across all samples.
@@ 2.1 Tully–Fisher samples
Our first method for obtaining redshift-independent distance estimates is the Tully–Fisher relation (Tully–Fisher relation; Tully_1977), which relates a spiral galaxy’s rotational velocity to its absolute magnitude $M$. Given an observed apparent magnitude, this relation yields a distance modulus and thereby a peculiar velocity, since the observed redshift is a function of the distance and peculiar velocity. The linewidth parameter $\eta$ is defined as
@@ 2.1.1 CosmicFlows-4 TFR sample
We use the CosmicFlows-4 Tully–Fisher relation catalogue, a subset of the broader CosmicFlows-4 compilation (Tully_2023), consisting of $9792$ galaxies with $z_{\rm CMB}\lesssim 0.05$ and no strict apparent magnitude limit (Kourkchi_2020B; Kourkchi_2020A). Our analysis uses photometry exclusively in the all-sky Wide-field Infrared Survey Explorer W1 band (same as Boubel_2025), with additional selection criteria applied. We require $\eta>-0.3$ (to eliminate dwarf and low-mass galaxies, which may follow a different Tully–Fisher relation or have a higher scatter), Galactic latitude $|b|>7.5^{\circ}$ to exclude the Galactic Zone of Avoidance, and a photometric quality flag of $5$ (“best”). After these cuts, the final W1 sample contains $3246$ galaxies. The highest photometric quality requirement excludes $\sim$$1300$ galaxies from the W1 subsample, while the Galactic latitude cut removes only 44 galaxies from the final sample. The Zone of Avoidance is typically defined as $|b|<5^{\circ}$ (e.g. Staveley-Smith_1998), making our choice relatively conservative. In contrast, Boubel_2025 do not apply either cut. Since the publicly released CosmicFlows-4 catalogue does not provide magnitude uncertainties, we adopt a conservative fiducial uncertainty of $0.05~\mathrm{mag}$, following Kourkchi_2019. However, this uncertainty is subdominant to the intrinsic scatter in the Tully–Fisher relation, which is approximately $0.35~\mathrm{mag}$ (see Fig.˜2 or Fig.˜9). Moreover, unlike Boubel_2025, who impose a lower redshift cut of $cz_{\rm CMB}>3000~\mathrm{km}\,\mathrm{s}^{-1}$, we do not apply a lower redshift limit in our main analysis, but discuss its impact in Section˜5.
@@
Next, we consider the 2MASS Tully–Fisher Survey (2MTF), an all-sky sample of 2,062 spiral galaxies with Tully–Fisher relation measurements extending to redshifts of $z_{\mathrm{CMB}}\lesssim 0.03$ (Masters_2008; Hong_2019). The survey is selected in the $K$ band with an apparent magnitude limit of $K<11.25$. We use the version of the catalogue compiled by Boruah_2019, which removes duplicates from the SFI++ sample, includes only $K$-band photometry, and applies a quality cut on linewidths, retaining only galaxies with $0.1<\eta<0.2$. These cuts leave $1247$ galaxies.
@@ 2.1.3 SFI++ TFR sample
We also consider the SFI++ catalogue, an all-sky sample comprising galaxies and groups with TFR measurements extending to redshifts of $z_{\mathrm{CMB}}\lesssim 0.05$ (Masters_2006; Springob_2007). Unlike 2MTF, the survey does not impose a strict apparent magnitude limit, and photometry is provided in the $I$ band. In this work, we use the galaxy-only version of the catalogue compiled by Boruah_2019, who apply quality cuts to select galaxies within the 2M++ footprint, impose a lower linewidth threshold to exclude low-mass galaxies, and use an iterative procedure to reject Tully–Fisher relation outliers. While Boruah_2019 also impose a strict upper linewidth selection to ensure linearity of the TFR, we relax this constraint and allow for Tully–Fisher relation curvature, resulting in a final sample of $2010$ galaxies.
@@ 2.2 Pantheon+ supernovae sample
Type Ia supernovae are widely used as standardisable candles in cosmology. The SALT2 model standardises their light curve (SALT2), yielding a standardised apparent magnitude via the Tripp formula (Tripp_1998):
@@
The Pantheon+ data set is a compilation of $1701$ spectroscopically confirmed Type Ia supernovae spanning redshifts from $z\sim 0.001$ to $\sim 2.3$ (Scolnic_2022; Brout_2022; Peterson_2022; Carr_2022). For consistency with our peculiar velocity modelling and to match the redshift range of the CosmicFlows-4 sample, we restrict the data to $z_{\rm CMB}\leq 0.05$, yielding a subset of $525$ supernovae. Pantheon+ combines the original Pantheon sample (Scolnic_2018) with updated low- and high-redshift supernovae, incorporating improved photometric calibration, light-curve standardisation, and systematic uncertainty modelling. Distance moduli are derived using the SALT2 light-curve fitter and corrected for selection effects via the BEAMS with Bias Corrections (BBC) method (Kessler_2017), which introduces an additive bias correction term in Eq.˜3. We use such bias-corrected apparent magnitudes $m_{\rm corr}$ from the Pantheon+ sample, which include a fiducial Tripp calibration. We therefore sample only $M_{\rm SN}$, not the light curve stretch and colour coefficients of the Tripp calibration.
@@
A reanalysis of the Pantheon+ data was presented by Lane_2024, who constructed a supernova covariance matrix designed to minimise dependence on the assumed cosmology and peculiar velocities. Unlike the standard Pantheon+ covariance matrix and bias-corrected apparent magnitudes, which assume a fiducial cosmology and Tripp parameters, the Lane_2024 covariance matrix does not include any contribution from the Tripp parametrisation. This enables simultaneous inference of the Tripp coefficients, with the standardised apparent magnitudes expressed as in Eq.˜3, and no further bias corrections. We treat this as a variant of the Pantheon+ data, denoted “Pantheon+Lane” henceforth.